OpenStudy (anonymous):

Cano went to Gator's Dockside for "all-you-can-eat" wings. The food bill totaled $25. He also left $2 for the tip. What percentage of the bill was his tip?

OpenStudy (anonymous):

\[\left( \frac{ x }{100 } \right)\left( 25 \right)=2\]

OpenStudy (anonymous):

what percent (x/100) of 25 is 2?

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

yes thank you.